The width of a rectangular lot is 20% of the length. If the perimeter is 192ft, then what are the length and width of the lot?
Found 2 solutions by solver91311, MrCrump:
Answer by solver91311(24713)



Two ways to look at this. The straight-forward way is to let represent the length. Then the width has to be .\

The perimeter is the sum of two times the length plus two times the width, so:



Just solve for

The other way is to realize that 20% is the same thing as one-fifth, so if we let represent the width, then must represent the length. Using the same perimeter formula:



Just solve for

John


Answer by MrCrump(1)

2x + 2(0.2x)=192
2x+.4x=192
2.4x=192
x=80
2(80)+2w=192
160+2w=192
2w=32
w=16
Length=80ft
Width=16ft